OpenStudy (anonymous):

how would i solve for a with this equation: -3a+3b/2 = -2a-c/-7

OpenStudy (anonymous):

did you forget parenthesis ? like (-3a+3b)/2 = (-2a-c)/-7 ? If not, a = 3b/2 - c/7

OpenStudy (anonymous):

you have three variables here. which one would you like to solve for?

OpenStudy (anonymous):

\[\frac{-3a+3b}{2}=\frac{-2a-c}{-7}\] like that one?

OpenStudy (gw2011):

If the problem is: -3a+(3b/2)=-2a-(c/-7), then the solution is the following: -3a+(3b/2)=-2a+(c/7) -3a+2a=-(3b/2)+(c/7) -a=-(3b/2)+(c/7) a=3b/2-c/7 Finding a common denominator gives you: a=(21b-2c)/14